M1=20kg
M2= 750kg
U1= 20m/s
U2= 15m/s
V1= 10m/s
V2= ?
1
Expert's answer
2019-09-30T10:02:57-0400

Solution:


According to the law of conservation of momentum:


"M_1\upsilon_1+M_2\upsilon_2=M_1u_1+M_2u_2"


"M_2\upsilon_2=M_1u_1+M_2u_2-M_1\upsilon_1"


Then:


"\upsilon_2=\frac{M_1u_1+M_2u_2-M_1\upsilon_1}{M_2}=\frac{M_1}{M_2}(u_1-\upsilon_1)+u_2"


"\upsilon_2=\frac{20}{750}(20-10)+15=15.27" m/s.


Answer:


"\upsilon_2" = 15.27 m/s.